Are you taking CLEP classes? Try clep.ai
| preferred AP College board partner for AP classes
medium Solved by 4 students

Bayes Theorem And Medical Testing

Mark
< Prev
Next >

Five percent of a population has a certain disease. A diagnostic test for the disease provides the results shown in the table, which lists the conditional probabilities of a test result given the patient’s disease status.

Test Result Disease Present Disease Absent
Test Positive 0.90 0.05
Test Negative 0.10 0.95

Consider the following statements:

I. The probability that an individual with the disease will test positive is 0.900.90.
II. The probability that an individual without the disease will test positive is 0.050.05.
III. The probability that an individual who tests positive actually has the disease is 0.05×0.050.05×0.05+0.95×0.90≈0.0029\frac{0.05 \times 0.05}{0.05 \times 0.05 + 0.95 \times 0.90} \approx 0.0029.

Which of the statements is/are true?

A

I only

B

I, II, and III

C

I and II only

D

II only

Hint
Did You Know?
Explain Why
Explain All Answers
Check Answer
Show Correct Answer
Report Question

Question Leaderboard

Rank
User
Correct Count
Attempt Count
Total Time
Score
#1minapup1711 0m 00s 100
#2halllais30411 1m 05s 35
#3abygail.gonzalez6912 1m 32s -2
#4282271428202 0m 00s -20
Items per page:
10
1 – 4 of 4

AI Tutor

How can I help?

APFIVE © 2020.
Email: apfive@apfive.org|Privacy Policy|